OpenStudy (anonymous):

Simplify the radical sqrt{5/9}

OpenStudy (precal):



OpenStudy (anonymous):

\[\LARGE\sqrt{\frac{5}{9}}=\frac{\sqrt5}{\sqrt9}=\frac{\sqrt5}{\sqrt{3^2}}=\frac{\sqrt5}{3}\]